$SPAD/src/input synonym.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1991-94.

———-

— * —

\getchunk{license}
-- This file contains the standard system defined system command
-- synonyms for the Axiom system. It is normally read into
-- the system by adding the line
--         )read synonym input )ifthere )quiet
-- to the start-up profile axiom.input.
-- If you wish to have a private list of synonyms, it is suggested
-- that you create a file mysyns.input.
-- This will also automatically be read into the system by the start-up
-- profile by the line
--         )read mysyns input )ifthere )quiet
-- The following synonyms are redefined in the system
-- )synonym  ?            what commands
-- )synonym  ap           what things
-- )synonym  apr          what things
-- )synonym  apropos      what things
-- )synonym  cache        set functions cache
-- )synonym  cl           clear
-- )synonym  cls          zsystemdevelopment )cls
-- )synonym  cms          system
-- )synonym  co           compiler
-- )synonym  d            display
-- )synonym  dep          display dependents
-- )synonym  dependents   display dependents
-- )synonym  disc         system cp disc
-- )synonym  e            edit
-- )synonym  expose       set expose add constructor
-- )synonym  fc           zsystemdevelopment )c
-- )synonym  fd           zsystemdevelopment )d
-- )synonym  fct          zsystemdevelopment )ct
-- )synonym  fctl         zsystemdevelopment )ctl
-- )synonym  fe           zsystemdevelopment )e
-- )synonym  fec          zsystemdevelopment )ec
-- )synonym  fect         zsystemdevelopment )ect
-- )synonym  fns          exec spadfn
-- )synonym  fortran      set output fortran

-- )synonym  h            help
-- )synonym  hd           system hypertex &
-- )synonym  kclam        boot clearClams ( )
-- )synonym  killcaches   boot clearConstructorAndLisplibCaches ( )
-- )synonym  patch        zsystemdevelopment )patch
-- )synonym  pause        zsystemdevelopment )pause
-- )synonym  prompt       set message prompt
-- )synonym  recurrence   set functions recurrence
-- )synonym  restore      history )restore
-- )synonym  save         history )save
-- )synonym  time         set message time
-- )synonym  type         set message type
-- )synonym  unexpose     set expose drop constructor
-- )synonym  up           zsystemdevelopment )update
-- )synonym  version      lisp *yearweek*
-- )synonym  w            what
-- )synonym  wc           what categories
-- )synonym  wd           what domains
-- )synonym  wp           what packages
-- )synonym  ws           what synonyms

———-

References^{[1] nothing}

5